You need to prepare an acetate buffer of pH 5.97 from a 0.724 M acetic acid solution and a 2.30 M KOH solution. If you have 575 mL of the acetic acid solution, how many milliliters of the KOH solution do you need to add to make a bufer of pH 5.97? The pKa of acetic acid is 4.76.
Answer #1

pH goal = 5.97

the buffer:

pH = pKa + log(A-/HA)

5.97 = 4.76 + log(A-/HA)

initially

mmol of HA = MV = 0.724*575 = 416.3 mmol of HA

mmol of A- = 0

after adding MV = 2.30*Vbase

mmol of HA = 416.3 -  2.30*Vbase

mmol of A- = 0+ 2.30*Vbase

5.97 = 4.76 + log(A-/HA)

5.97 = 4.76 + log( ( 2.30*Vbase) /(416.3 -  2.30*Vbase))

10^(5.97 - 4.76) = ( 2.30*Vbase) /(416.3 -  2.30*Vbase)

16.218 * (416.3-  2.30*Vbase) =  2.30*Vbase

16.218 * (416.3) - 16.218 *2.3*Vbase = 2.3*Vbase

Vvase (  16.218 *2.3+2.3) = 16.218 * (416.3)

Vbase = 16.218 * (416.3) / (  16.218 *2.3+2.3)

Vbase= 170.487 mL required